OpenStudy (anonymous):

write this equation in point slope form and slope intercept form ( 2,-7) ( 3,5)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(2\quad ,&-7)\quad &(3\quad ,&5) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ \textit{point-slope form}\implies y-y_1=m(x-x_1)\\ \quad \\ \textit{slope-point intercept form}\implies y=mx+b\)

OpenStudy (anonymous):

IM THE MAN U NEED

OpenStudy (isaiah.feynman):

Lol.

OpenStudy (jdoe0001):

sounds like the title of an R&B song =P

OpenStudy (jdoe0001):

\(\bf \implies y-y_1=m(x-x_1)\qquad \textit{plug in your values}\\ \quad \\ \implies y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\\\textit{ for the slope intercept form}\)

OpenStudy (anonymous):